LLP is auditing Dumbledore Manufacturing, Inc. (DMI) for the year ended December 31, 2018. You are the staff auditor responsible for preparing the financial statements. With the unadjusted trial balance provided, you are responsible for the following:

Recording any adjusting journal entries based on audit findings on both the trial balance and the AJE report. Don’t forget to add a description.

  1. When auditing other assets, it was determined that DMI did not amortize goodwill in 2018. Goodwill is amortized over 10 years. DMI realized that in order to keep the account on track for the 10 year amortization they needed to record an entry to correct this in 2018.
  1. When auditing accrued expenses, it was discovered that DMI failed to accrue for the last pay period of December wages. DMI’s pay schedule is bi-weekly (10 working days) and the payroll report for that full 2-week period leading up to the end of the year totaled $241,957. Per review of the calendar, 6 days should have been accrued for at December 31, 2018. Payroll expenses are allocated in the following manner:
    1. COGS – 45%
    2. Salary – 40%
    3. Hourly – 15%

Due to the magnitude of the error, DMI needs to record an adjusting entry to correct. Round your entry to the nearest dollar (no pennies).
Answer #1

a) Journal Entry for goodwill amortization

Amortization of goodwill a/c Dr. (Goodwill amount /10)

To Goodwill a/c Cr. (Goodwill amount/10)

(Being goodwill amortization recognized recorded for the year as per useful life of 10 years)

b) COGS A/c Dr. $ 119,769

Salary expense A/c Dr. $ 106,641

Wages A/c Dr. $ 39,923  

To Accruals A/c Cr. $2,66,153

(Being accrual of expense recorded in relation to payroll for last 2 weeks of December 31, 2018 in respective expense heads)

Calculation of amount to be accrued

Total payroll expense for 10 days= $241,957

Proportionate amount for 11 days (considering 6 days in last week)= ($241,957/10*11)= $2,66,153

Distribution of amount to relevant expense head

COGS= $ 266,153*45%= $ 119,769

Salary= $ 266,153*40%= $ 106,641

Wages= $ 266,153*15%= $ 39,923